# Unifying Attention Heads and Task Vectors via Hidden State Geometry in In-Context Learning

### v1

@@ Abstract
The unusual properties of in-context learning (ICL) have prompted investigations into the internal mechanisms of large language models. Prior work typically focuses on either special attention heads or task vectors at specific layers, but lacks a unified framework linking these components to the evolution of hidden states across layers that ultimately produce the model’s output. In this paper, we propose such a framework for ICL in classification tasks by analyzing two geometric factors that govern performance: the separability and alignment of query hidden states. A fine-grained analysis of layer-wise dynamics reveals a striking two-stage mechanism—separability emerges in early layers, while alignment develops in later layers. Ablation studies further show that Previous Token Heads drive separability, while Induction Heads and task vectors enhance alignment. Our findings thus bridge the gap between attention heads and task vectors, offering a unified account of ICL’s underlying mechanisms.
@@ 1 Introduction
One of the most remarkable features of Large Language Models (LLMs) is their ability to respond to novel queries in user-desired mannerisms on-the-fly solely by learning from demonstrations provided in the input (as shown in Figure 1 (A))—without any additional training. This capability is known as In-context Learning (ICL) [Brown et al., 2020, Radford et al., 2019]. ICL has revolutionized natural language processing by reducing dependence on burdensome data collection and costly finetuning, enabling swift and seamless adaptation of LLMs to a variety of downstream tasks [Dong et al., 2024].
@@
Therefore, in this work, we propose a framework as shown in Figure 1 for ICL in classification tasks by a key geometric insight: the classification performance of an LLM depends on whether the hidden states of queries with different ground-truth labels are (1) sufficiently separable, and (2) well aligned with the unembedding vectors of labels. In other words, hidden states should be separated along a direction that aligns with the difference between the unembedding vectors of different labels. To investigate this, we collect hidden states of queries in both ICL and zero-shot settings, and analyze them using a range of geometric measures to assess how ICL improves hidden states in terms of separability and alignment. Our approach operates on three levels: (1) We examine the layer-wise trends in these measures and compare zero-shot with ICL (and across different ICL settings) to quantify the effect of ICL on the hidden states geometry. (2) We observe how changes in measure values reflect the evolution of semantic information represented in the hidden states that ultimately informs model outputs in ICL. (3) We probe specific layers to examine the role of different types of attention heads and how they influence both the values and trajectories of these geometric measures.
Our experiments find that ICL improves classification performance primarily by enhancing alignment between query hidden states and unembedding vectors, with less impact on separability. Our results further suggest that ICL proceeds as a two-stage process: early layers primarily increase the separability of hidden states, while in middle-to-late layers, separability plateaus and alignment improve. The refinement of alignment is achieved through layer updates that filter out semantic information irrelevant to the task and preserve label-relevant content, thereby semantically refining the query hidden states. Such updates qualify ICL query hidden states as task vectors.
To build the bridge between the task vector and model components, we analyze the effect of attention heads on these geometric measures and find that (as demonstrated by Figure 1): (1) The early increase in separability is driven by a proliferation of Previous Token Heads (PTHs)—attention heads that attend to the previous token at each position—in early layers. (2) In middle-to-late layers, Induction Heads (IHs) enhance alignment by amplifying hidden state components along the unembedding directions of the labels. Thus, our findings not only explain why attention heads like PTHs and IHs are essential for ICL performance, but also clarify why induction head outputs empirically serve as effective task vectors [Todd et al., 2024, Hendel et al., 2023].
@@ 2 Related works
IH and PTH   Previous works on the mechanistic cause of ICL observe structural circuits in Transformers composed of a Previous Token Head (PTH) and an Induction Head (IH), where the PTH collects information at demonstration label tokens, and the IH copies it to query positions for prediction [Olsson et al., 2022, Singh et al., 2024, Song et al., 2025]. Their importance in classification has been observed in large-scale models as well [Cho et al., 2025a]. However, existing analyses focus on end-to-end effects and leave open how these heads influence query hidden states for classification.
Task vectors   An alternative theory proposes that LLMs summarize demonstrations into task vectors, which can be used to steer zero-shot hidden states to achieve ICL-level prediction accuracy. These vectors may be ICL query hidden states extracted at an intermediate layer [Hendel et al., 2023, Liu et al., 2024] or attention head outputs [Todd et al., 2024, Li et al., 2024]. Decoding task vectors through the unembedding matrix reveals that they encode task-related tokens and support semantic manipulations akin to “word2vec”-style algebra [Mikolov et al., 2013]. Despite intriguing empirical findings, explanations for their mechanism remain vague.
@@ 3 Theoretical framework
A classification task involves prompting an LLM with a query $x$ and ground-truth label $y$ (assumed to be a single token), e.g., $x$ = “I like this movie. Sentiment:” and $y$ = “positive”. The LLM with $L$ layers encodes $x$ into a vector $\bm{h}_{x}\in\mathbb{R}^{d}$ taken from the hidden state at the final token (“:”). Given a vocabulary ${\mathbb{V}}=\{v_{1},...,v_{|{\mathbb{V}}|}\}$ and output embedding matrix $\bm{E}\in\mathbb{R}^{|{\mathbb{V}}|\times d}$, the model predicts a distribution: $p(v_{j}|x)={\exp(\bm{E}_{v_{j}}\bm{h}_{x})}/{\sum_{j^{\prime}=1}^{|{\mathbb{V}}|}\exp(\bm{E}_{v_{j^{\prime}}}\bm{h}_{x})},$ where $\bm{E}_{v_{j}}$ is the unembedding vector of $v_{j}$. The model then outputs the token $v_{j}$ with the highest probability as the label for $x$.
In the ICL scenario, as shown in Figure 1 (A), $k$ demonstration-label pairs $(x_{1},y_{1}),...,(x_{k},y_{k})$ are prepended to form the prompt $[x_{1},y_{1},...,x_{k},y_{k},x]$, which modifies the hidden state of $x$ and results in a different distribution $p(v_{j}|x_{1},y_{1},...,x_{k},y_{k},x)$, potentially improving prediction accuracy.
@@
where the first equality states that classification accuracy equals the proportion of the labels being predicted as the most probable token and the second equality uses the monotonicity of softmax.
We first consider the simple case where $\{x_{1},...,x_{n}\}$ only has $2$ different labels, $y_{A}$ and $y_{B}$, i.e. $y_{i}\in\{y_{A},y_{B}\},\forall i$, and leave the generalization to the case of more labels to Appendix B. For a unit vector $\bm{u}\in{\mathbb{R}}^{d}$, we define the separability of $\bm{H}$ along direction $\bm{u}$ with labels $y_{A},y_{B}$ to be:
@@
where ${\mathbb{N}}_{A}=\{i:y_{i}=y_{A}\}$ and ${\mathbb{N}}_{B}=\{i:y_{i}=y_{B}\}$. $S(\bm{u})$ measures the fraction of hidden states that can be correctly classified by a linear decision boundary orthogonal to $\bm{u}$. By considering all possible directions, we obtain the maximum separability of $\bm{H}$, i.e. $S^{*}=\sup_{\bm{u}\in{\mathbb{R}}^{d},\|\bm{u}\|_{2}=1}S(\bm{u})$. Though $S(\bm{u})$ is not a continuous function, $S^{*}$ as a supremum can be attained on ${\mathbb{S}}^{d-1}=\{\bm{u}:\bm{u}\in{\mathbb{R}}^{d},\|\bm{u}\|_{2}=1\}$ (See Appendix A). Denote $\bm{u}^{*}$ as the $\bm{u}$ where $S$ attains the supremum, then:
@@ 4 Method: Measuring separability and alignment of hidden states
Given the aforementioned influence of ICL, attention heads, and task vectors on the separability and alignment of hidden states across layers, we introduce several measures [^2] For binary classification datasets, each measure is computed over the entire $\bm{H}$. For multiclass datasets with labels $y_{1},...,y_{m}$, we enumerate all label pairs $(y_{j},y_{k})$, compute the measure using the corresponding two label clusters, i.e. hidden states of queries with labels $y_{j},y_{k}$, before averaging across all pairs to obtain a scalar summary for $\bm{H}$. As shown in Appendix B, this is justified by the fact that separability and alignment influence the accuracy of multiclass classification through the binary classification between each label pair. to quantify their influence. The formal definitions and mathematical details of all measures are provided in Appendix C.
@@ A. Separability measure: Separability score
Theorem 1 shows that the classification is affected by the maximum separability of hidden states. However, finding maximum separability requires evaluating separability along infinitely many directions in ${\mathbb{S}}^{d-1}$ and is intractable. As a practical proxy, we train a logistic classifier on a subset of $\bm{H}$ and evaluate its accuracy on held-out data. We call the accuracy separability score as it reflects the empirical maximum separability of hidden states achieved by the classifier.
@@ B. Alignment measures
In Theorem 1 we show that both output alignment and directional alignment affect accuracy. Below, we introduce measures for both.
@@ 5 Experiments
Datasets   We conduct experiments on two major types of classification datasets: text classification (SUBJ [Wang and Manning, 2012], SST-2 [Socher et al., 2013], TREC [Li and Roth, 2002]) and natural language inference (SNLI [MacCartney and Manning, 2008], RTE [Dagan et al., 2005], and CB [De Marneffe et al., 2019]). We also use a generated dataset to explore the scenario where ground-truth labels of queries are not presented in the demonstrations (discussed in Subsection H.2).
ICL setting   We set the number of in-context demonstrations to be 8 unless otherwise stated. The demonstration samples are chosen randomly, except for one experiment that investigates the effect of using demonstrations procured by kNN retrieval [Liu et al., 2022]. For a detailed exposition of the implementation of models, datasets, prompt templates, etc., refer to Appendix D.
@@ 5.1 Layer-wise trends of separability and alignment
ICL as a two-stage process   Figure 2 shows the values of six geometric measures at all layers of Llama2-70B in both zero-shot and ICL settings averaged across datasets (output alignment is deferred to Subsection 5.2 due to its strong link to the semantic information of hidden states; results for other models are in Appendix E). A key observation is that Figure 2 reveals a striking phase transition in the evolution along the layers of ICL hidden states. Phase 1 (Figure 1 (B)): In the early layers, directional alignment measures take low values, whereas separability increases rapidly. Phase 2 (Figure 1 (D)): In the subsequent layers, this trend reverses: separability plateaus, while all directional alignment measures spike simultaneously, which suggests that the label unembedding difference direction starts to capture more variance and separate label clusters more effectively. It causes the semantic information of query labels to be predominant in $\bm{H}$, which ultimately leads to final output alignment. These findings are consistent with prior work showing that task information emerges only after the initial layers [Sia et al., 2024]. The alternation between two phases is also captured by the effective dimension of $\bm{H}$, which first rises and then declines as alignment starts to improve, reflecting a gradual concentration of hidden states along label unembedding directions. In contrast, this transition is much weaker in the zero-shot setting, where effective dimension rises monotonously and the rise in alignment measures is not nearly as pronounced. Hence, this phase transition pattern characterizes ICL as a two-stage process: an initial phase enhancing separability, followed by a refinement phase that improves alignment.
@@
ICL improves accuracy primarily through alignment   Another key observation is the small gap in separability scores between ICL and zero-shot settings—despite an 80% (Figure 2, upper middle) difference in classification accuracy. This suggests that the semantics-based separation of query hidden states is an inherent property of LLM inference, not a major result of ICL, and that alignment is the primary bottleneck in the zero-shot case. As visualized in Figure 1 (C), the direction of high separability in zero-shot hidden states is misaligned with the label-difference vector, as indicated by low singular and composite alignment scores. Along this direction, label clusters remain poorly separated (low mean-based alignment) and account for little variance (low variance-based alignment). These findings align with prior work showing that unembedding vectors often fail to reflect the main axes of separation in hidden state space [Cho et al., 2025b].
A Varying number of demonstrations
B KNN demonstration selection
C Symbolized demonstration labels
@@
Given the volatility of ICL accuracy across settings with different demonstration numbers and input formats, we investigate whether the phase transition persists across settings, and whether accuracy differences can similarly be attributed to alignment properties of hidden states. We consider three settings: (1) Varying the number of demonstrations from 0 to 24 in increments of 4; (2) Changing the demonstration selection strategy to kNN retrieval where demonstrations with closest zero-shot embeddings to each query are chosen; (3) Replacing original demonstration labels with semantically uninformative symbols like @ and #. Figure 3 reports the results averaged across datasets.
Phase transition is a robust hallmark of ICL   Figure 3 shows that the phase transition pattern persists across all three settings. In Figure 3 (A), it is evident at all demonstration counts except 0. Figure 3 (B) confirms that principled demonstration selection (via kNN) preserves instead of interfering with the transition. Even in Figure 3 (C), where labels are replaced by symbols, the two-stage pattern in effective dimension remains—distinguishing it from zero-shot, where effective dimension rises monotonically—though the post-transition alignment surge is dampened.
Alignment explains accuracy differences across ICL settings   Accuracy differences between ICL settings are consistently reflected in alignment measures. In Figure 3 (A), increasing the number of demonstrations from 4 to 24 produces minimal changes in separability, while alignment measures improve in a consistent manner as the trend curves are shifted versions of one another. Effective dimension also decreases steadily, suggesting a higher concentration of hidden states along label-unembedding directions. Figure 3 (B) similarly shows that the accuracy gains from kNN-selected demonstrations correspond to stronger alignment and lower effective dimension. In Figure 3 (C), the accuracy drop from symbolic label replacement aligns with a flattened post-transition rise in alignment measures, resembling the case of zero-shot.

### v2

@@ Abstract
The unusual properties of in-context learning (ICL) have prompted investigations into the internal mechanisms of large language models. Prior work typically focuses on either special attention heads or task vectors at specific layers, but lacks a unified framework linking these components to the evolution of hidden states across layers that ultimately produce the model’s output. In this paper, we propose such a framework for ICL primarily in classification tasks by analyzing two geometric factors that govern performance: the separability and alignment of query hidden states. A fine-grained analysis of layer-wise dynamics reveals a striking two-stage mechanism—separability emerges in early layers, while alignment develops in later layers. Ablation studies further show that Previous Token Heads drive separability, while Induction Heads and task vectors enhance alignment. Our findings thus bridge the gap between attention heads and task vectors, offering a unified account of ICL’s underlying mechanisms. [^1] Code implementation: https://github.com/HLYang2001/ICL_Hidden_Geometry.
@@ 1 Introduction
One of the most remarkable features of Large Language Models (LLMs) is their ability to respond to novel queries in user-desired mannerisms on-the-fly solely by learning from demonstrations provided in the input (as shown in Figure˜1 (A))—without any additional training. This capability is known as In-context Learning (ICL) [Brown et al., 2020, Radford et al., 2019]. ICL has revolutionized natural language processing by reducing dependence on burdensome data collection and costly finetuning, enabling swift and seamless adaptation of LLMs to a variety of downstream tasks [Dong et al., 2024].
@@
Therefore, in this work, we propose a framework, shown in Figure˜1, based on a key geometric insight into LLM classification: the success of LLMs in classification depends on whether the hidden states of queries with different ground-truth labels are (1) sufficiently separable and (2) well aligned with the unembedding vectors of labels. In other words, hidden states should be separated along a direction that aligns with the difference between the unembedding vectors of different labels. To investigate this, we collect hidden states in both ICL and zero-shot settings and analyze them using a range of geometric measures to assess how ICL improves hidden states in terms of separability and alignment. Our approach operates on three levels: (1) We examine the layer-wise trends in these measures and compare zero-shot with ICL (and across different ICL settings) to quantify the effect of ICL on the hidden state geometry. (2) We observe how changes in measure values reflect the evolution of semantic information represented in the hidden states that ultimately inform model outputs in ICL. (3) We probe specific layers to examine the role of different types of attention heads and how they influence both the values and trajectories of these geometric measures.
Our experiments find that ICL improves classification performance primarily by enhancing alignment between query hidden states and unembedding vectors, with less impact on separability. Results further suggest that ICL proceeds as a two-stage process: early layers mainly increase hidden-state separability, while in middle-to-late layers, separability plateaus and alignment improves. The refinement of alignment arises through layer updates that filter out task-irrelevant semantics and preserve label-relevant content, thereby semantically refining query hidden states. Such updates qualify ICL query hidden states as task vectors. We also test this two-stage characterization on ICL for a generation task and confirm its explanatory power across settings.
To build the bridge between the task vector and model components, we analyze the effect of attention heads on these geometric measures and find that (as demonstrated by Figure˜1): (1) The early increase in separability is driven by a proliferation of Previous Token Heads (PTHs)—attention heads that attend to the previous token at each position—in early layers. (2) In middle-to-late layers, Induction Heads (IHs) enhance alignment by amplifying hidden state components along the unembedding directions of the labels. Thus, our findings not only explain why attention heads like PTHs and IHs are essential for ICL performance, but also clarify why induction head outputs empirically serve as effective task vectors [Todd et al., 2024, Hendel et al., 2023].
@@ 2 Related works
IH and PTH  Previous works on the mechanistic cause of ICL identify structural circuits in Transformers comprising a Previous Token Head (PTH) and an Induction Head (IH), where the PTH gathers information at demonstration label tokens, and the IH copies it to query positions for prediction [Olsson et al., 2022, Singh et al., 2024, Song et al., 2025]. Their importance in classification has been confirmed in large-scale models as well [Cho et al., 2025a]. However, existing analyses focus on end-to-end effects, leaving open how these heads shape query hidden states.
Task vectors  An alternative theory proposes that LLMs summarize demonstrations into task vectors, which can be used to steer zero-shot hidden states to increase prediction accuracy. These vectors may be ICL query hidden states extracted at an intermediate layer [Hendel et al., 2023, Liu et al., 2024], attention head outputs [Todd et al., 2024, Li et al., 2024], or MLP outputs [Merullo et al., 2024]. Decoding task vectors through the unembedding matrix reveals that they encode task-related tokens and support semantic manipulations akin to “word2vec”-style algebra [Mikolov et al., 2013]. Despite intriguing empirical findings, explanations for their mechanism remain preliminary. Kahardipraja et al. [2025] show that heads expressing contextual information or model’s parametric knowledge produce task vectors with distinct function. Jiang et al. [2025b] and Han et al. [2025] study how the geometry of model’s task representation impacts the effects of injecting task vectors. Jiang et al. [2025a] reveal the correlation between model’s catastrophic forgetting and the change in task vectors. However, none of these studies answer the question of how task vectors influence the concrete internal computation process of models which directly and ultimately determine the outputs.
@@ 3 Theoretical framework
We begin our framework with an analysis of classification tasks, where an LLM is prompted with a query $x$ and its ground-truth label $y$, e.g., $x=$ “I like this movie. Sentiment:” and $y=$ “positive”. The LLM with $L$ layers encodes $x$ into a vector $\bm{h}_{x}\in\mathbb{R}^{d}$, taken from the hidden state at the final token (“:”). Given a vocabulary ${\mathbb{V}}=\{v_{1},...,v_{|{\mathbb{V}}|}\}$ and an output embedding matrix $\bm{E}\in\mathbb{R}^{|{\mathbb{V}}|\times d}$, the model predicts $p(v_{j}|x)={\exp(\bm{E}_{v_{j}}\bm{h}_{x})}/{\sum_{j^{\prime}=1}^{|{\mathbb{V}}|}\exp(\bm{E}_{v_{j^{\prime}}}\bm{h}_{x})},$ where $\bm{E}_{v_{j}}$ is the unembedding vector of $v_{j}$. The model outputs the token $v_{j}$ with the highest probability as the predicted label for $x$.
In the ICL scenario, as shown in Figure˜1 (A), $k$ demonstration-label pairs $(x_{1},y_{1}),...,(x_{k},y_{k})$ are prepended to form the prompt $x_{1},y_{1},...,x_{k},y_{k},x]$, which modifies the hidden state of $x$ and results in a different distribution $p(v_{j}|x_{1},y_{1},...,x_{k},y_{k},x)$, potentially improving prediction accuracy.
@@
where the first equality states that classification accuracy equals the proportion of the labels being predicted as the most probable token, and the second equality uses the monotonicity of softmax.
We first consider the simple case where $\{x_{1},...,x_{n}\}$ only has $2$ different labels, $y_{A}$ and $y_{B}$, i.e. $y_{i}\in\{y_{A},y_{B}\},\forall i$, and leave the generalization to the case of more labels to Appendix˜B. For a unit vector $\bm{u}\in{\mathbb{R}}^{d}$, we define the separability of $\bm{H}$ along direction $\bm{u}$ with labels $y_{A},y_{B}$ to be:
@@
where ${\mathbb{N}}_{A}=\{i:y_{i}=y_{A}\}$ and ${\mathbb{N}}_{B}=\{i:y_{i}=y_{B}\}$. $S(\bm{u})$ measures the fraction of hidden states that can be correctly classified by a linear decision boundary orthogonal to $\bm{u}$. By considering all possible directions, we obtain the maximum separability of $\bm{H}$, i.e. $S^{*}=\sup_{\bm{u}\in{\mathbb{R}}^{d},\|\bm{u}\|_{2}=1}S(\bm{u})$. Though $S(\bm{u})$ is not a continuous function, $S^{*}$ as a supremum can be attained on ${\mathbb{S}}^{d-1}=\{\bm{u}:\bm{u}\in{\mathbb{R}}^{d},\|\bm{u}\|_{2}=1\}$ (See Appendix˜A). Denote $\bm{u}^{*}$ as the $\bm{u}$ where $S$ attains the supremum, then:
@@ 4 Method: Measuring separability and alignment of hidden states
Given the aforementioned influence of ICL, attention heads, and task vectors on the separability and alignment of hidden states across layers, we introduce several measures [^3] For binary classification datasets, each measure is computed over the entire $\bm{H}$. For multiclass datasets with labels $y_{1},...,y_{m}$, we enumerate all label pairs $(y_{j},y_{k})$, compute the measure using the corresponding two label clusters, i.e. hidden states of queries with labels $y_{j},y_{k}$, before averaging across all pairs to obtain a scalar summary for $\bm{H}$. As shown in Appendix B, this is justified by the fact that separability and alignment influence the accuracy of multiclass classification through the binary classification between each label pair. to quantify their influence. The formal definitions and mathematical details of all measures are provided in Appendix˜C.
@@ A. Separability measure: Separability score
Theorem˜1 shows that the classification is affected by the maximum separability of hidden states. However, finding maximum separability requires evaluating separability along infinitely many directions in ${\mathbb{S}}^{d-1}$ and is intractable. As a practical proxy, we train a logistic classifier on a subset of $\bm{H}$ and evaluate its accuracy on held-out data. We call the accuracy separability score as it reflects the empirical maximum separability of hidden states achieved by the classifier.
@@ B. Alignment measures
In Theorem˜1 we show that both output alignment and directional alignment affect accuracy. Below, we introduce measures for both.
@@ 5 Experiments and Results
Datasets  We conduct experiments on two major types of classification datasets: text classification (SUBJ [Wang and Manning, 2012], SST-2 [Socher et al., 2013], TREC [Li and Roth, 2002]) and natural language inference (SNLI [MacCartney and Manning, 2008], RTE [Dagan et al., 2005], and CB [De Marneffe et al., 2019]). We further include a generation dataset (detailed in Subsection˜5.1) to test the generalizability of our results in settings other than classification. We also use a generated dataset to explore the scenario where ground-truth labels of queries are not presented in the demonstrations (discussed in Subsection˜I.5).
ICL setting  We set the number of in-context demonstrations to be 8 unless otherwise stated. The demonstration samples are chosen randomly, except for one experiment that investigates the effect of using demonstrations procured by kNN retrieval [Liu et al., 2022]. For a detailed exposition of the implementation of models, datasets, prompt templates, etc., refer to Appendix˜D.
@@ 5.1 Layer-wise trends of separability and alignment
ICL as a two-stage processFigure˜2 shows the values of six geometric measures at all layers of Llama2-70B in both zero-shot and ICL settings averaged across datasets (output alignment is deferred to Subsection˜5.2 due to its strong link to the semantic information of hidden states; results for other models are in Appendix˜F). A key observation is that Figure˜2 reveals a striking phase transition in the evolution along the layers of ICL hidden states. Phase 1 (Figure˜1 (B)): In the early layers, directional alignment measures take low values, whereas separability increases rapidly. Phase 2 (Figure˜1 (D)): In the subsequent layers, this trend reverses: separability plateaus, while all directional alignment measures spike simultaneously, which suggests that the label unembedding difference direction starts to capture more variance and separate label clusters more effectively. It causes the semantic information of query labels to be predominant in $\bm{H}$, which ultimately leads to final output alignment. These findings are consistent with prior work showing that task information emerges only after the initial layers [Sia et al., 2024]. The alternation between two phases is also captured by the effective dimension of $\bm{H}$, which first rises and then declines as alignment starts to improve, reflecting a gradual concentration of hidden states along label unembedding directions. In contrast, this transition is much weaker in the zero-shot setting, where effective dimension rises monotonously and the rise in alignment measures is not nearly as pronounced. Hence, this phase transition pattern characterizes ICL as a two-stage process: an initial phase enhancing separability, followed by a refinement phase that improves alignment.
@@
ICL improves accuracy primarily through alignment  A key observation is that ICL and zero-shot settings show only a small gap in separability scores—despite an 80% (Figure˜2, upper middle) difference in classification accuracy. This implies that the semantic separation of query hidden states is intrinsic to LLM inference rather than a main effect of ICL (see Subsection˜E.1 for the practical implication of this), making alignment the primary bottleneck in zero-shot performance. As shown in Figure˜1 (C), the high-separability direction in zero-shot hidden states is misaligned with the label-difference vector, as reflected in low singular and composite alignment scores. Along this direction, label clusters remain poorly separated (low mean-based alignment) and explain little variance (low variance-based alignment). These findings align with prior work indicating that unembedding vectors often fail to capture the dominant separation axes in hidden state space [Cho et al., 2025b].
Given the volatility of ICL accuracy across settings with different demonstration numbers and input formats, we investigate whether the phase transition persists across settings, and whether accuracy differences can similarly be attributed to alignment properties of hidden states. We consider three settings: (1) Varying the number of demonstrations from 0 to 24 in increments of 4; (2) Changing the demonstration selection strategy to kNN retrieval where demonstrations with the closest zero-shot embeddings to each query are chosen; (3) Replacing original demonstration labels with semantically uninformative symbols like “@” and “#”. Figure˜3 reports the results averaged across datasets.
(a) Varying number of demonstrations
(b) KNN demonstration selection
(c) Symbolized demonstration labels
@@
Phase transition is a robust hallmark of ICLFigure˜3 shows that the phase transition pattern persists across all three settings. In Figure˜3 (A), it is evident at all demonstration counts except 0. Figure˜3 (B) confirms that principled demonstration selection (via kNN) preserves instead of interfering with the transition. Even in Figure˜3 (C), where labels are replaced by symbols, the two-stage pattern in effective dimension remains—distinguishing it from zero-shot, where effective dimension rises monotonically—though the post-transition alignment surge is dampened.
Phase transition holds in open-ended settings  To test whether the phase transition pattern occurs in tasks beyond classification, i.e., open-ended generative tasks, where ICL is also frequently applied, we consider the following task: the queries assume the format “Praise/Critique {subject}”, where {subject} is the name of a food (e.g., tapas). The labels are reviews for the subject satisfying the sentiment requirement, generated using GPT-4o [OpenAI et al., 2024], for instance, “Praise tapas $\rightarrow$ The tapas assortment at this restaurant is absolutely delightful!”. We track the layer-wise dynamics of the separability score and composite alignment [^4] Since a generation task has an indefinite and infinite label space, we let the label unembedding difference direction needed to calculate composite alignment be that between “positive” and “negative”. as representatives of the metrics evaluated on the 8-shot and 0-shot hidden states of this task. The results in Figure 4 mirror those in Figure 2, where the few-shot and zero-shot hidden states exhibit similar separability properties across layers, but the surge in composite alignment beginning from mid-layers is specific to few-shot hidden states.
(a) Layer-wise separability score: 0-shot vs. 8-shot
(b) Layer-wise composite alignment: 0-shot vs. 8-shot
Figure 4: (A) In the generation setting, the separability score exhibits almost identical trends in the 0-shot and 8-shot cases. (B) In the 8-shot case, the phase-transition of composite alignment is pronounced.
Alignment explains accuracy differences across ICL settings  We find that: accuracy differences between ICL settings are consistently reflected in alignment measures. Specifically, in Figure˜3 (A), increasing the number of demonstrations from 4 to 24 produces minimal changes in separability, while alignment measures improve in a consistent manner as the trend curves are shifted versions of one another. Effective dimension also decreases steadily, suggesting a higher concentration of hidden states along label-unembedding directions. Figure˜3 (B) similarly shows that the accuracy gains from kNN-selected demonstrations correspond to stronger alignment and lower effective dimension. In Figure˜3 (C), the accuracy drop from symbolic label replacement aligns with a flattened post-transition rise in alignment measures, resembling the case of zero-shot.